OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = x^2 – 1 y = –x – 1

OpenStudy (anonymous):

y = x^2 – 1 y = –x – 1 -x - 1 = x^2 - 1 x^2 + x = 0 x(x + 1) = 0 x = 0, -1 y = 0^2 – 1 y = -1 y = -1^2 – 1 y = 0

sam (.sam.):

y = x^2 – 1 y = –x – 1 ------------------------- x^2 – 1 = –x – 1 x^(2)+x-1=-1 x(x+1)=0 x=0 x+1=0 x=-1 x=0,-1 ================= y = –x – 1 y=-1, y=0